# Ultra-low value for the allowed decay of 110Agm confirmed via mass measurements

### v1

@@ Abstract
The mass of 109Ag isotope was measured with respect to 110Cd with the phase-imaging ion-cyclotron-resonance technique using the JYFLTRAP double Penning trap at the IGISOL facility. By combining this new result with the known spectroscopic information, the value of eV between the isomer of 110Ag and the state in 110Cd was determined. This is the lowest for any allowed transition observed to date. The nuclear shell-model results were performed with the kshell code employing the jj45pnb interaction, and combined with the state-of-the-art atomic calculations. The theoretical partial half-life, years, and the resulting branching ratio, , combined with a viable production method, thermal-neutron capture on stable 109Ag, make 110Agm as a promising candidate for future antineutrino mass measurement experiments.
The determination of the absolute mass of neutrinos remains one of the big unanswered questions in physics. Currently, the most stringent limits come from the kinematic experiments: KATRIN [1, 2] and MARE [3, 4] for the electron antineutrino, and ECHo [5, 6] and HOLMES [7, 8] for the electron neutrino. They are based on the model-independent analysis of the end-point spectrum of decays of 3H (KATRIN) and 187Re (MARE) or 163Ho electron capture (ECHo and HOLMES). Another tritium-based experiment, Project 8, is currently being planned [9]. These three isotopes have been chosen due to the very small ground-state-to-ground-state values ( eV for 3H [10], eV for 163Ho [11] and eV for 187Re [12]). Nevertheless, despite such narrow energy windows, only a very small fraction of decays ends up near the endpoint, limiting the sensitivity of these experiments.
An alternative approach has been recently proposed where the decay proceeds to the excited state in the daughter isotope with an ultra-low value, usually keV [13]. So far, the most promising candidate is 115In where the second-forbidden unique decay to the first excited states in 115Sn () was experimentally observed [14, 15] and its value is around 200 eV [16, 14]. Two other isotopes of interest are 135Cs where keV for the first-forbidden unique decay to the excited state [17] as well as 131I with the value of keV for an allowed decay to the excited state. [18]. For both 135Cs and 131I, these ultra-low- decays are still awaiting experimental confirmation.
@@
One of isomers which might be of interest for the future antineutrino mass experiments is 110Agm ( keV, , d [32, 33]). It decays almost exclusively (98.67(8)%) via to the excited states in stable 110Cd (see Fig. 1) [32, 33] and it can be easily produced in nuclear reactors thanks to the large cross section ( b) for the thermal-neutron capture on stable 109Ag [34, 35]. The literature value is 3008.3(13) keV [36, 32] and the recent studies constrained the excitation energy of the state to keV [37], resulting in the present energy window for the allowed iso-ex decay of keV. Additionally, the state is not isomeric as it deexcites via prompt rays [38, 39]. Consequently, 110Agm is a perfect candidate for a potential ultra-low- transition.
@@
The masses of 109Ag and 110Ag isotopes are connected via a precise single-neutron separation energy value from the study ( keV [40]) while the 110Agm and excitation energies are known from -ray spectroscopy studies [33]. As a result, can be expressed as a function of masses of two stable isotopes, 109Ag and 110Cd, and precise spectroscopic observables. Considering that the quadratically-added uncertainties of , and are equal to 127 eV, the keV uncertainty of [36, 32] can be reduced if the single-proton separation energy () of 110Cd is known with a precision better than 100 eV. In order to accomplish it, the mass of 109Ag was determined with 110Cd as a reference by utilizing the JYFLTRAP double Penning trap mass spectrometer [41] at the Ion Guide Isotope Separator On-Line (IGISOL) facility at the University of Jyväskylä [42]. Additionally, to check for possible mass-dependent systematic effects caused by the u mass difference, two measurements using zirconium oxide molecules, 90Zr16O and 92Zr16O against 91Zr16O, were also conducted.
@@
The silver ions were produced by a discharge ion source placed in the target area and extracted using a sextupole ion guide [43] while the ions of cadmium and zirconium were produced by an offline discharge ion source [44]. After extraction, the ions were accelerated by a 30 kV electrostatic potential and transported through a 55∘ dipole magnet. Then, the continuous beam was sent into the gas-filled radio-frequency quadrupole cooler-buncher [45]. In there, in addition to cooling and bunching the beam, the zirconium oxide molecules were created. Subsequently, the bunches were injected into the first trap of JYFLTRAP, where the axial motion of the bunch was damped by collisions with the gas and mass-selective buffer-gas cooling technique [46] applied to select the ions of interest.
@@
In the second trap, the phase-imaging ion-cyclotron-resonance (PI-ICR) technique [47, 48, 49, 50] was used to measure the ions’ free space cyclotron frequency , where and are the charge and mass of the ion and the magnetic field strength. Two timing patterns were used to excite the ions’ radial eigenmotions. For the magnetron motion, the pattern started with a dipolar radiofrequency pulse exciting the reduced cyclotron motion. This was followed by a quadrupolar pulse converting the motion to magnetron motion after an accumulation time . Finally, the ions were projected onto a position-sensitive microchannel plate detector (2D-MCP). For the reduced cyclotron motion, the order of the converting pulse and the accumulation time was reversed. Typical phase spots obtained in this work are shown in Fig. 2a) and b).
The cyclotron frequency was determined using the angle between the phase spots after magnetron () and cyclotron () motion accumulation, and the number of revolutions the ions performed during :
The mass difference of an ion-of-interest and a reference was determined from the ratio of their measured frequencies.
To reduce the effect of residual magnetron and cyclotron motion on the ions’ position, the start of the conversion pulse was scanned over the cyclotron period and the extraction from the measurement trap over the period of the magnetron motion. The effect of the temporal magnetic field fluctuation was minimized by alternating the measurements of the ion of interest and the reference ion every seven minutes as the fluctuation of the JYFLTRAP magnetic field has been measured to be min-1 [50]. By performing count-rate class analysis [51, 52, 50] using bunches with up to five detected ions, the effect of ion-ion interaction was determined to be negligible.
@@
For the mass determination of 109Ag, a total of 157 frequency ratios between 110Cd and 109Ag were measured over 40 hours, see Fig. 2c). To account for other sources of uncertainties, the data were collected using three different accumulation times, 813 ms, 1 s and 1.5 s, and two different spot positions on the detector. However, no significant effects were observed based on these factors. A final ratio was obtained as a weighted average of all the measurements. Similarly, for 92Zr16O mass determination, 98 ratios over 25 hours were measured using two different spot positions and two accumulation times of 813 ms and 1 s, while for 90Zr16O, 20 ratios with s were collected over five hours, both against the 91Zr16O reference, see Supplemental Material [53].
@@
The updated keV and the mass excess of 109Ag (ME = keV) agree with AME20 ( keV, ME = keV [36]) and they are 26 and three times more precise, respectively, with the uncertainty of the latter being dominated by the uncertainty of the 110Cd mass. These results allows for reevaluation of the value:
where is the frequency ratio between 110Cd and 109Ag from this work, u is the atomic mass of 110Cd [36], and are the masses of an electron and a neutron, respectively, keV [32] and keV [40] are the 110Agm excitation energy and the 110Ag single-neutron separation energy, respectively, while eV and eV are the first ionization energies of Cd and Ag, respectively [54].
@@
All the -decay energies extracted in this work agree with AME20 [36] but they are an order of magnitude more precise, see Table 2. The updated is positive, eV, at level. It is the lowest observed value for an allowed transition and the second-lowest after the second-forbidden unique decay of 115In to the first excited state in 115Sn [31].
To estimate the partial half-lives for the decays of 110Ag to the excited states in 110Cd, we utilized the kshell code [58] within the Nuclear Shell Model (NSM) framework, employing the jj45pnb interaction [59], a two-nucleon potential based on a perturbative G-matrix approach. The model space includes the proton orbitals , , , and with single-particle energies of , , , and MeV, respectively. For neutrons, the model space comprises the , , , , and orbitals, with corresponding single-particle energies of , , , , and MeV, as specified in the jj45pnb interaction. The computations were performed without truncation of the model space. We note that for the applied region of , only variations of the jj45pnb interaction are available, yielding similar (or redundant) results within the available Hamiltonians found in codes such as kshell or NuShellX@MSU [60].
| Transition | (keV) | ||||||
| Exp. | 0.5 | 0.6 | 0.7 | 0.8 | |||
| (d) | 808.51 | 561.61 | 390.007 | 286.536 | 219.379 | 0.0117 | 528.88 |
| (d) | 369.02 | 117.627 | 81.6857 | 60.0139 | 45.9482 | -0.4742 | 82.07 |
| (s) | 25.58 | 21.85 | 15.18 | 11.15 | 8.54 | 0.7271 | 2891.22 |
| ( y) | 111This transition has not been observed experimentally yet. | 2.2309 | 1.5492 | 1.1382 | 0.8714 | 0.1542 | 0.405 |
To validate the applied interaction, we computed the excitation energies, magnetic dipole and electric quadrupole moments for both 110Ag and 110Cd, and compared them with evaluated data, see Supplemental Material [53]. The results demonstrated reasonable agreement, except for the quadrupole moment of the 110Ag state. We note that we used standard effective charges: for protons and for neutrons, and the bare g-factors , , , and , and did not attempt to adjust them to reproduce the electromagnetic moments values accurately.
@@
Since the phase-space factor is extremely sensitive to the decay energy, for the transition the expected half-lives for the different values and for the energies ranging from 225 to 600 eV are plotted, see Fig. 3. The experimental uncertainties found in this work are indicated by a gray shaded region. The partial half-life ranges from y for the upper limit (540 eV) with up to y for the lower limit (270 eV) with . For any given curve, we mostly observe an order of magnitude difference in whilst considering the gray shaded area. The partial half-lives for two values, 0.5 and 0.8, were also computed without incorporating the atomic exchange corrections [61] to assess its impact. As can be seen in Fig. 3, the values increase by a factor , and they range between and years. This result highlights the importance of this correction for the ultra-low values.
@@
In summary, we have measured with high precision the single-proton separation energy of 110Cd, keV, resulting in the improved mass-excess value of 109Ag, keV. By combining our results with the well known single-neutron separation energy and the isomer excitation energy of 110Ag, we demonstrated that the -decay transition has the ultra-low value of 405(135) eV. This is the lowest for any allowed transition measured to date and the first low case measured for the -decaying isomer. The nuclear shell model combined with the atomic calculations predicted the partial half-life for this transition to be years which translates to the branching ratio of . Consequently, our work indicates that 110Agm might be a good candidate for the future antineutrino mass measurement experiments. It also demonstrates that the future search of low- transitions should not be limited to the -decaying ground states, with the isomers in 101,102Rh, 106Ag, 120Sb, 121Te, 184Re and 194Ir being the most promising cases to find such transitions.
In passing we note that the hints of the decay have been already observed. In two 110Agm decay studies, the 846-keV ray, which is the strongest transition deexciting the state in 110Cd [37], was detected [63, 64]. In addition, a second transition deexciting this level, the 1466-keV ray, was observed in one of these studies [63]. Both of them remain unplaced [63, 64, 33]. However, it should be noted that if these rays originate from the deexcitation of the state, the partial decay half-life to this state would be in the years range, which is orders of magnitude shorter than indicated by calculations. Thus, a future dedicated study to search for the decay is needed.
@@ Acknowledgements.
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 771036 (ERC CoG MAIDEN), from the European Union’s Horizon Europe Research and Innovation Programme under Grant Agreement No. 101057511 (EURO-LABS) and the Research Council of Finland projects No. 295207, 306980, 327629, 354589 and 354968. J.R. acknowledges financial support from the Vilho, Yrjö and Kalle Väisälä Foundation.References
@@
- Ge et al. [2021a] Z. Ge, T. Eronen, A. de Roubin, D. A. Nesterenko, M. Hukkanen, O. Beliuskina, R. de Groote, S. Geldhof, W. Gins, A. Kankainen, A. Koszorús, J. Kotila, J. Kostensalo, I. D. Moore, A. Raggio, S. Rinta-Antila, J. Suhonen, V. Virtanen, A. P. Weaver, A. Zadvornaya, and A. Jokinen, Direct measurement of the mass difference of rules out as a promising -decay candidate to determine the neutrino mass, Phys. Rev. C 103, 065502 (2021a).
- Ge et al. [2021b] Z. Ge, T. Eronen, K. S. Tyrin, J. Kotila, J. Kostensalo, D. A. Nesterenko, O. Beliuskina, R. de Groote, A. de Roubin, S. Geldhof, W. Gins, M. Hukkanen, A. Jokinen, A. Kankainen, A. Koszorús, M. I. Krivoruchenko, S. Kujanpää, I. D. Moore, A. Raggio, S. Rinta-Antila, J. Suhonen, V. Virtanen, A. P. Weaver, and A. Zadvornaya, Electron-Capture: A New Candidate for Neutrino Mass Determination, Phys. Rev. Lett. 127, 272301 (2021b).
@@
- Ge et al. [2024b] Z. Ge, T. Eronen, V. A. Sevestrean, O. Niţescu, S. Stoica, M. Ramalho, J. Suhonen, A. de Roubin, D. Nesterenko, A. Kankainen, P. Ascher, S. A. S. Andres, O. Beliuskina, P. Delahaye, M. Flayol, M. Gerbaux, S. Grévy, M. Hukkanen, A. Jaries, A. Jokinen, A. Husson, D. Kahl, J. Kostensalo, J. Kotila, I. Moore, S. Nikas, M. Stryjczyk, and V. Virtanen, High-precision measurements of the atomic mass and electron-capture decay value of 95Tc (2024b), arXiv:2406.05273 [nucl-ex] .
@@
- Karvonen et al. [2008] P. Karvonen, I.D. Moore, T. Sonoda, T. Kessler, H. Penttilä, K. Peräjärvi, P. Ronkanen, and J. Äystö, A sextupole ion beam guide to improve the efficiency and beam quality at IGISOL, Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms 266, 4794 (2008).

### v2

@@ Abstract
The mass of the electron-antineutrino can be determined in dedicated measurements of the spectral shape near the endpoint of a transition, with a low value enhancing the sensitivity of the measurement. One such low--value candidate is the transition between the isomer of 110Ag and the state in 110Cd ( keV). To reduce the uncertainty of the value, we have used the phase-imaging ion-cyclotron-resonance technique with the JYFLTRAP double Penning trap and performed a high-precision atomic-mass measurement of 109Ag with 110Cd as a reference. Combined with the known spectroscopic data, we obtain a re-evaluated value eV, for the transition. This represents the lowest value for any allowed transition observed to date. In order to estimate the partial half-life () and branching ratio (Br) of the transition, nuclear shell-model calculations were performed using the Hamiltonian in combination with state-of-the-art atomic calculations. The computed values years and , along with the thermal-neutron capture on stable 109Ag as a viable production method, make a promising candidate for future antineutrino-mass measurements.
The determination of the absolute mass of neutrinos remains one of the big unanswered questions in physics. Currently, the most stringent limits come from the kinematic experiments: KATRIN [1, 2] and MARE [3, 4] for the electron antineutrino, and ECHo [5, 6] and HOLMES [7, 8] for the electron neutrino. They are based on the model-independent analysis of the end-point spectrum of decays of 3H (KATRIN) and 187Re (MARE) or 163Ho electron capture (ECHo and HOLMES). Another tritium-based experiment, Project 8, is currently being planned [9]. These three isotopes have been chosen due to the very small ground-state-to-ground-state values ( eV for 3H [10], eV for 163Ho [11] and eV for 187Re [12]). Nevertheless, despite such narrow energy windows, only a very small fraction of decays ends up near the endpoint, limiting the sensitivity of these experiments. In neutrino mass experiments involving decay and electron capture (EC), a low value enhances sensitivity to (anti)neutrino mass. Thus, finding such decays with the smallest possible value is desirable for experiments aimed at determining the (anti)neutrino mass.
An alternative approach has been recently proposed where the decay proceeds to the excited nuclear state in the daughter isotope with an ultra-low value, usually keV [13]. The first promising -decay candidate discovered is 115In where the second-forbidden unique decay to the first excited states in 115Sn () was experimentally observed [14, 15] and its value is around 200 eV [16, 14]. Two other -decay isotopes of interest are 135Cs where keV for the first-forbidden unique decay to the excited state [17] as well as 131I with the value of keV for an allowed decay to the excited state. [18]. Additionally, a promising EC candidate 159Dy was recently discovered [19] with the of 1.18(19) keV for the 159Dy(3/2-) 159Tb∗(5/2-) transition. However, with the exception of 115In, these ultra-low- decays are still awaiting experimental confirmation.
@@
One isomer that could be of interest for future antineutrino mass experiments is 110Agm ( keV, , d [32, 33]). It decays predominantly (98.67(8)%) via to the excited states in stable 110Cd (see Fig. 1), with a minor decay branch (1.33(8)%) via internal transition [32, 33]. Additionally, it can be readily produced in nuclear reactors due to the large cross section ( b) for thermal-neutron capture on stable 109Ag [34, 35]. The literature value is 3008.3(13) keV [36, 32] and the recent studies constrained the excitation energy of the non-isomeric state to keV [37], resulting in the present energy window for the allowed iso-ex decay of keV. Interestingly, hints of the state being populated in the -decay of 110Agm have already been observed. The 846 keV ray, which is the strongest transition de-exciting the state [37] was detected in two studies [38, 39]. In addition, a second transition de-exciting this level, the 1466 keV ray, was observed in one of these studies [38]. However, both of them remain unplaced in the -decay of 110Ag [38, 39, 33]. All these facts make 110Agm a perfect candidate for a potential ultra-low- transition.
@@
The masses of 109Ag and 110Ag isotopes are connected via a precise single-neutron separation energy value from the study ( keV [40]) while the 110Agm and excitation energies are known from -ray spectroscopy studies [33]. As a result, can be expressed as a function of masses of two stable isotopes, 109Ag and 110Cd, and precise spectroscopic observables. Considering that the quadratically-added uncertainties of , and are equal to 127 eV, the keV uncertainty of [36, 32] can be reduced if the single-proton separation energy () of 110Cd is known with a precision better than 100 eV. In order to accomplish it, the mass of 109Ag was determined with 110Cd as a reference by utilizing the JYFLTRAP double Penning trap mass spectrometer [41] at the Ion Guide Isotope Separator On-Line (IGISOL) facility at the University of Jyväskylä [42]. This measurement results in the most precise determination of the , which is presented in this letter. To evaluate the potential of this transition for neutrino mass detection, we incorporated the precise value into nuclear shell-model calculations using the KSHELL code with the jj45pnb interaction, combined with advanced atomic calculations. Additionally, to check for possible mass-dependent systematic effects caused by the u mass difference, two mass measurements of zirconium oxide molecules, 90Zr16O and 92Zr16O with 91Zr16O as a reference, were also conducted, see Supplemental Material [43]. These molecules are in the same mass region as 109Ag with nuclear masses that are well characterized [44, 45, 36, 32].
@@
To produce the ions used in the measurements, a sample of natural silver was placed in an electric discharge ion source located in the target area of the IGISOL [42], while samples of natural cadmium and zirconium were placed in the offline discharge ion source station [46]. The extracted ions were accelerated by a 30 kV electrostatic potential. The ions were transported through a 55∘ dipole magnet capable of separating the isobars containing the desired ions, 109Ag+, 110Cd+ and 90-92Zr+, from the ion beam and allowing them to enter the radio-frequency quadrupole cooler-buncher [47]. In addition to cooling and bunching the beam for more efficient injection into the JYFLTRAP, the ZrO+ molecules were formed there. Subsequently, the bunches were injected into the helium-filled preparation trap of JYFLTRAP [41], where the mass-selective buffer-gas cooling technique [48] was applied to separate the ions of interest from possible isobaric contaminants. The ions of interest were then sent to the measurement trap of JYFLTRAP.
In the measurement trap, the phase-imaging ion-cyclotron-resonance (PI-ICR) technique [49, 50, 51, 52] was used to measure the free space cyclotron frequency of an ion species, where and are the charge and mass of the ion and the magnetic field strength. By measuring the cyclotron frequencies of an ion of interest () and a reference ion (), we obtain the ratio of the mass of the ion of interest to the mass of the reference from the ratio of their frequencies, as
With the PI-ICR technique, is determined by projecting the ions in the trap onto a position-sensitive microchannel plate detector (2D-MCP). The cyclotron frequency is obtained using the angle between the centers of the spots formed on the detector relative to the center spot after magnetron () and reduced cyclotron () eigenmotion accumulation
@@
where is the accumation time and the number of revolutions during at frequency . Two timing patterns were used to excite the ions’ radial eigenmotions using radiofrequency electric fields in order to produce the spots. At the start of both patterns, a dipolar pulse with frequency was used to reduce the initial magnetron motion of the ions to bring them to the center of the trap. Then for the magnetron spot, the pattern continued with a dipolar radiofrequency pulse with frequency exciting the reduced cyclotron motion. This was followed by a quadrupolar pulse with frequency converting the motion to magnetron motion. Finally, after waiting for an accumulation time , the ions were projected onto the 2D-MCP registering their final positions. For the reduced cyclotron spot, the order of the converting pulse and the accumulation time was reversed. The position of the center of the trap on the 2D-MCP was determined by extracting the centered ions without further excitations. Typical position data obtained in this work are shown in Fig. 2a) and b).
@@
For the mass determination of 109Ag, a total of 157 frequency ratios between 109Ag+ and the reference 110Cd+ were measured over 40 hours using three different accumulation times, 813 ms, 1 s and 1.5 s, see Fig. 2c). Two different spot positions on the detector were used to account for distortion in the projected image, but no systematic shift in the measured ratio was observed. A final ratio was obtained as a weighted average of all the measurements. Similarly for 92Zr16O+ mass determination, 98 ratios over 25 hours were measured using two different spot positions and two accumulation times of 813 ms and 1 s, while for 90Zr16O+, 20 ratios with s were collected over five hours, both against the 91Zr16O+ reference, see Supplemental Material [43].
To reduce the effect of residual magnetron and cyclotron motion on the ions’ position, the start of the conversion pulse was scanned over the cyclotron period ( 1 s) and the extraction from the measurement trap over the period of the magnetron motion ( 600 s). The effect of the temporal magnetic field fluctuation was minimized by alternating the measurements of the ion of interest and the reference ion every seven minutes and linearly interpolating the frequency of the reference at the time of the ion of interest measurement. The temporal fluctuation of the JYFLTRAP magnetic field has been measured to be min-1 [52]. By performing count-rate class analysis [53, 54, 52] using bunches with up to five detected ions, the effect of ion-ion interactions was determined to be negligible.
@@
The updated proton separation energy, keV, and the mass excess of 109Ag (ME = keV) are consistent with the values reported in AME20 ( keV, ME = keV [36]). These updated values are 26 and three times more precise, respectively, with the uncertainty in the latter being dominated by the uncertainty in the mass of 110Cd. These results allows for reevaluation of the value:
where is the frequency ratio between 110Cd+ and 109Ag+ from this work, u is the atomic mass of 110Cd [36], and are the masses of an electron and a neutron, respectively, keV [32] and keV [40] are the 110Agm excitation energy and the 110Ag single-neutron separation energy, respectively, while eV and eV are the first ionization energies of Cd and Ag, respectively [55].
@@
All the -decay energies extracted in this work agree with AME20 [36] but they are an order of magnitude more precise, see Table 2. The updated value confirms that the transition is energetically allowed at a confidence level () of , with eV. It is the lowest observed value for an allowed transition and the second-lowest overall after the second-forbidden unique decay of 115In to the first excited state in 115Sn [31].
To estimate the partial half-lives for the decays of 110Ag to the excited states in 110Cd, we utilized the kshell code [57] within the Nuclear Shell Model (NSM) framework, employing the jj45pnb interaction [58], a two-nucleon potential based on a perturbative G-matrix approach. The model space includes the proton orbitals , , , and with single-particle energies of , , , and MeV, respectively. For neutrons, the model space comprises the , , , , and orbitals, with corresponding single-particle energies of , , , , and MeV, as specified in the jj45pnb interaction. The computations were performed without truncation of the model space. We note that for the nuclear mass region around , only variants of the Hamiltonian are available in the libraries of the shell-model codes KSHELL and NuShellX@MSU [59]. All these variants produce similar results so that we consider here the results obtained by using the Hamiltonian as representative of all these Hamiltonians.
| Transition | (keV) | |||||||
| Exp. | 0.5 | 0.6 | 0.7 | 0.8 | Units | |||
| 561.61 | 390.007 | 286.536 | 219.379 | d | 0.0117 | 528.88(13) | ||
| 117.627 | 81.6857 | 60.0139 | 45.9482 | d | -0.4742 | 82.07(13) | ||
| 111This transition has not been observed experimentally yet. | 22.309 | 15.492 | 11.382 | 8.714 | My | 0.1542 | 0.405(135) | |
| 21.85 | 15.18 | 11.15 | 8.54 | s | 0.7271 | 2891.22(12) | ||
To validate the applied interaction, we computed the excitation energies, magnetic dipole and electric quadrupole moments for both 110Ag and 110Cd, and compared them with evaluated data, see Supplemental Material [43]. In computations of the electromagnetic moments, we have used standard effective charges for protons and for neutrons, and the bare g-factors , , , and . This choice leads to a reasonable agreement between the computed and measured electromagnetic moments, indicating that the wave functions of the involved nuclear states are fairly reliable. The computed excitation energies of the and states, involved in the ultra-low- transition, correspond rather well to their experimental counterparts, thus speaking for their reliability in the half-life and branching estimations.
@@
Since the phase-space factor is extremely sensitive to the decay energy, for the transition the expected half-lives for the different values and for the energies ranging from 225 to 600 eV are plotted, see Fig. 3. The experimental uncertainties found in this work are indicated by a gray shaded region. The partial half-life ranges from y for the upper limit (540 eV) with up to y for the lower limit (270 eV) with . For any given curve, we mostly observe an order of magnitude difference in whilst considering the gray shaded area. The partial half-lives for two values, 0.5 and 0.8, were also computed without incorporating the atomic exchange corrections [60] to assess its impact. As can be seen in Fig. 3, the values increase by a factor , and they range between and years. This result highlights the importance of this correction and emphasizes the importance of high-precision measurements of the ultra-low for the half-life determination. The observed discrepancies between theoretical predictions and experimental half-lives can be linked to limitations of the Nuclear Shell Model. These include the exclusion of orbitals outside the defined model space, the neglect of nuclear deformation effects, and the lack of fine-tuning of interaction parameters for this isotope. While addressing these issues would require significant effort, the results still serve as a valuable benchmark for evaluating the current computational framework.
@@
In summary, we have measured with high precision the single-proton separation energy of 110Cd, keV, resulting in the improved mass-excess value of 109Ag, keV. By combining our results with the well known single-neutron separation energy and the isomer excitation energy of 110Ag, we demonstrated that the -decay transition has the ultra-low value of 405(135) eV. This is the lowest for any allowed transition measured to date and the first low case measured for a -decaying isomer. The nuclear shell model combined with the atomic calculations predicted the partial half-life for this transition to be years which translates to the branching ratio of . It should be noted that if the 846 keV and 1466 keV rays observed in the decay studies of 110Ag originate from the de-excitation of the state in 110Cd, the partial decay half-life to this state would be in the years range, which is orders of magnitude shorter than indicated by calculations. All in all, our work indicates that 110Agm might be a good candidate for the future antineutrino mass measurement experiments, although a new decay study of 110Agm is required to confirm the production of the state in 110Cd. Our work also demonstrates that the future search of low- transitions should not be limited to the -decaying ground states, with the isomers in 101,102Rh, 106Ag, 120Sb, 121Te, 184Re and 194Ir being the most promising cases to find such transitions.
@@ Acknowledgements.
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 771036 (ERC CoG MAIDEN), from the European Union’s Horizon Europe Research and Innovation Programme under Grant Agreement No. 101057511 (EURO-LABS) and the Research Council of Finland projects No. 295207, 306980, 327629, 354589 and 354968. J.R. acknowledges the financial support from the Vilho, Yrjö and Kalle Väisälä Foundation.References
@@
- Ge et al. [2021a] Z. Ge, T. Eronen, K. S. Tyrin, J. Kotila, J. Kostensalo, D. A. Nesterenko, O. Beliuskina, R. de Groote, A. de Roubin, S. Geldhof, W. Gins, M. Hukkanen, A. Jokinen, A. Kankainen, A. Koszorús, M. I. Krivoruchenko, S. Kujanpää, I. D. Moore, A. Raggio, S. Rinta-Antila, J. Suhonen, V. Virtanen, A. P. Weaver, and A. Zadvornaya, Electron-Capture: A New Candidate for Neutrino Mass Determination, Phys. Rev. Lett. 127, 272301 (2021a).
@@
- Ge et al. [2021b] Z. Ge, T. Eronen, A. de Roubin, D. A. Nesterenko, M. Hukkanen, O. Beliuskina, R. de Groote, S. Geldhof, W. Gins, A. Kankainen, A. Koszorús, J. Kotila, J. Kostensalo, I. D. Moore, A. Raggio, S. Rinta-Antila, J. Suhonen, V. Virtanen, A. P. Weaver, A. Zadvornaya, and A. Jokinen, Direct measurement of the mass difference of rules out as a promising -decay candidate to determine the neutrino mass, Phys. Rev. C 103, 065502 (2021b).
@@
- Ge et al. [2024b] Z. Ge, T. Eronen, V. A. Sevestrean, O. Niţescu, S. Stoica, M. Ramalho, J. Suhonen, A. de Roubin, D. Nesterenko, A. Kankainen, P. Ascher, S. Ayet San Andres, O. Beliuskina, P. Delahaye, M. Flayol, M. Gerbaux, S. Grévy, M. Hukkanen, A. Jaries, A. Jokinen, A. Husson, D. Kahl, J. Kostensalo, J. Kotila, I. Moore, S. Nikas, M. Stryjczyk, and V. Virtanen, High-precision measurements of the atomic mass and electron-capture decay value of 95Tc, Physics Letters B 859, 139094 (2024b).